OpenStudy (anonymous):

what is the slope of (2,0) (0,-6)

OpenStudy (anonymous):

formula for finding slope (m)= (y2 -y1)/(x2-x1)

OpenStudy (anonymous):

in your case slope=(-6-0)/(0-2)= (-6)/(-2)=3